# Exact and efficient basis pursuit denoising via differential inclusions and a selection principle

### v1

@@ Appendix B Detailed Proofs
[…]
By assumption that $t_{0}>0$ and continuity, there is some $\epsilon_{0}>0$ such that $0\leqslant t_{0}-\epsilon_{0}$, $t\in[t_{0}-\epsilon_{0},t_{0}]$, and $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$ remains the solution to problem (70). It will remain the same as $\epsilon_{0}$ decreases until either $\epsilon_{0}=t_{0}$ or there exists $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ with $|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|\neq\boldsymbol{0}$ such that $\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{q})=-|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|/(1-t/t_{0})$, that is, a constraint becomes satisfied with equality. In the latter case, we can rearrange this expression to obtain
$t=t_{0}\left(1-\frac{|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|}{|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{q})|}\right).$
The first index $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ with $\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})\neq 0$ and $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\leqslant-|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|$, if it exists, is the one whose ratio $|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|/|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{q})|$ is minimized. Hence
$T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})\coloneqq t_{0}\left(1-\inf_{\begin{subarray}{c}j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\\ \boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})\neq 0\\ \hat{\boldsymbol{v}}_{j}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\leqslant-|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|\end{subarray}}\frac{|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|}{|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{q})|}\right)$
is the smallest number lesser than $t_{0}$ for which $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$ solves problem (70).
Now, consider the descent direction (69), suppose $t\in[\max(0,T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})),t_{0}]$ and let
$\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q})$
@@
$\boldsymbol{d}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=(1-t/t_{0})\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}).$
The descent direction depends linearly on $t$, and so the corresponding maximal descent time $\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ in Proposition 3.1(i) is inversely proportional to $t$. More precisely:
$\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})/(1-t/t_{0}).$ (71)
@@
$C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq$
$\qquad\min_{j\in\{1,\dots,n\}}\left\{\frac{\mathrm{sgn}{\left(\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),\boldsymbol{e}_{j}\rangle\right)}-\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle)}{\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),\boldsymbol{e}_{j}\rangle}\right\}.$
@@
$\boldsymbol{d}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\Delta(1-t/t_{0})\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q});t,\boldsymbol{b}-(t-t_{0})\boldsymbol{q})=(1-t\Delta)(1-t/t_{0})\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}).$
We now seek the smallest nonnegative number $T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})\leqslant t_{0}$ in terms of $\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ for which
$1-t\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=\boldsymbol{0}$
for every $t\in[T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q}),t_{0}]$. Equation (71) gives
$1-t\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=\boldsymbol{0}$ $\iff 1-tC^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})/(1-t/t_{0})=0$
$\iff 1-t/t_{0}+tC^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})=0$
$\iff t=t_{0}/(1+t_{0}C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})).$
@@
$T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq\frac{t_{0}}{1+t_{0}C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})}.$
@@
$t_{1}\coloneqq\max(T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s}),T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})),$
for every $t\in(t_{1},t_{0}]$ there exists some $\Delta\in[0,\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}))$ for which $01-t\Delta=$ 0. Note that $0\leqslant t_{1}<t_{0}$ since $0\leqslant T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})<t_{0}$ and $T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})<t_{0}$. In particular, we have that $t_{1}=0\implies T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})=0$ and
$T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})=0\iff C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})=+\infty\iff\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})=\boldsymbol{0}.$
@@
$\boldsymbol{d}\left(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}\right)=\boldsymbol{0}.$
@@
$\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=\begin{cases}&\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\quad\text{if $t_{1}>0$},\\ &\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})\quad\text{otherwise,}\end{cases}$
is the solution to (dBPDN) at hyperparameter $t$ and data $\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}$ on $[t_{1},t_{0}]$. Furthermore,
$t\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ $=\left(\frac{t}{t_{0}}\right)t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$
@@
$\qquad\qquad+\left(1-\frac{t}{t_{0}}\right)(\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q}))$
$=t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\left(\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})+t_{0}\boldsymbol{q}\right)$
$=\bm{A}\left[\boldsymbol{x}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\right]-\boldsymbol{b}-(t_{0}-t)\boldsymbol{q},$

### v2

@@ Appendix B Technical Proofs
$\text{$\boldsymbol{v}_{j}\geqslant-\infty$ if $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ and $|\boldsymbol{x}^{s}(t_{0},\boldsymbol{b})|\neq\boldsymbol{0}$}.$
By assumption that $t_{0}>0$ and continuity, there is some $\epsilon_{0}>0$ such that $0\leqslant t_{0}-\epsilon_{0}$, $t\in[t_{0}-\epsilon_{0},t_{0}]$, and $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b})$ remains the solution to problem (57). It will remain the same as $\epsilon_{0}$ decreases until either $\epsilon_{0}=t_{0}$ or there exists $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ with $|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|\neq\boldsymbol{0}$ such that $\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b})=-|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|/(1-t/t_{0})$, that is, a constraint becomes satisfied with equality. In the latter case, we can rearrange this expression to obtain
$t=t_{0}\left(1-\frac{|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|}{|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b})|}\right).$
The first index $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ with $\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})\neq 0$ and $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b})\leqslant-|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|$, if it exists, is the one whose ratio $|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|/|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b})|$ is minimized. Hence
$T_{-}(t_{0},\boldsymbol{b},\hat{\boldsymbol{v}}^{s})\coloneqq t_{0}\left(1-\inf_{\begin{subarray}{c}j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\\ \boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})\neq 0\\ \hat{\boldsymbol{v}}_{j}^{s}(t_{0},\boldsymbol{b})\leqslant-|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|\end{subarray}}\frac{|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|}{|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b})|}\right)$
is the smallest number lesser than $t_{0}$ for which $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b})$ solves problem (57).
Now, consider the descent direction (56), suppose $t\in[\max(0,T_{-}(t_{0},\boldsymbol{b},\hat{\boldsymbol{v}}^{s})),t_{0}]$ and let
$\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})\coloneqq\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b})+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$
@@
$\boldsymbol{d}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b})=(1-t/t_{0})\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b}).$
The descent direction depends linearly on $t$, and so the corresponding maximal descent time $\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b})$ in Proposition 4.1(i) is inversely proportional to $t$. More precisely:
$\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b})=C^{s}(t_{0},\boldsymbol{b})/(1-t/t_{0}).$ (58)
@@
$C^{s}(t_{0},\boldsymbol{b})\coloneqq\min_{j\in\{1,\dots,n\}}\left\{\frac{\mathrm{sgn}{\left(\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle\right)}-\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle)}{\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle}\right\}.$
@@
$\boldsymbol{d}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\Delta(1-t/t_{0})\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b})=(1-t\Delta)(1-t/t_{0})\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b}).$
We now seek the smallest nonnegative number $T_{+}(t_{0},\boldsymbol{b})\leqslant t_{0}$ in terms of $\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b})$ for which
$1-t\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b})=\boldsymbol{0}$
for every $t\in[T_{+}(t_{0},\boldsymbol{b}),t_{0}]$. Equation (58) gives
$1-t\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b})=\boldsymbol{0}$ $\iff 1-tC^{s}(t_{0},\boldsymbol{b})/(1-t/t_{0})=0$
$\iff 1-t/t_{0}+tC^{s}(t_{0},\boldsymbol{b})=0$
$\iff t=t_{0}/(1+t_{0}C^{s}(t_{0},\boldsymbol{b})).$
@@
$T_{+}(t_{0},\boldsymbol{b})\coloneqq\frac{t_{0}}{1+t_{0}C^{s}(t_{0},\boldsymbol{b})}.$
@@
$t_{1}\coloneqq\max(T_{-}(t_{0},\boldsymbol{b},\hat{\boldsymbol{v}}^{s}),T_{+}(t_{0},\boldsymbol{b})),$
for every $t\in(t_{1},t_{0}]$ there exists some $\Delta\in[0,\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}))$ for which $1-t\Delta=0$. Note that $0\leqslant t_{1}<t_{0}$ since $0\leqslant T_{+}(t_{0},\boldsymbol{b})<t_{0}$ and $T_{-}(t_{0},\boldsymbol{b},\hat{\boldsymbol{v}}^{s})<t_{0}$. In particular, we have that $t_{1}=0\implies T_{+}(t_{0},\boldsymbol{b})=0$ and
$T_{+}(t_{0},\boldsymbol{b})=0\iff C^{s}(t_{0},\boldsymbol{b})=+\infty\iff\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})=\boldsymbol{0}.$
@@
$\boldsymbol{d}\left(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}\right)=\boldsymbol{0}.$
@@
$\boldsymbol{p}^{s}(t,\boldsymbol{b})=\begin{cases}&\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})\quad\text{if $t_{1}>0$},\\ &\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})\quad\text{otherwise,}\end{cases}$
is the solution to (dLASSO) at hyperparameter $t$ and data $\boldsymbol{b}$ on $t_{1},t_{0}]$. Furthermore,
$t\boldsymbol{p}^{s}(t,\boldsymbol{b})$ $=\left(\frac{t}{t_{0}}\right)t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})$
@@
$\qquad\qquad+\left(1-\frac{t}{t_{0}}\right)(\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})))$
$=t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\left(\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b})\right)$
$=\bm{A}\left[\boldsymbol{x}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b})\right]-\boldsymbol{b},$